Can you help me with this assignment in analytic geometry:

Show that each of the points are the vertices of a rectangle
L ( -4,3); M (0,-2); N(5,2); O (1,7)
I need a complete solution. Can you graph it. Thank you!

Answer by Boreal(15235)

Slope of LM=-5/4, point slope of line is y-y1=m(x-x1), m slope, (x1,y1) point
y+2=(-5/4)x, and y=-(5/4)x-2
-----------------------
MN needs a slope of 4/5, to be perpendicular to LM, and the slope is 4/5
y+2=(4/5)(x) and y=(4/5)x+2
-----------------------
NO needs a slope of -5/4, and and it is
y-2=-(5/4)(x-5)
y=-(5/4)x+33/4
-----------------------
OL needs a slope of 4/5, and it is
y-7=(4/5)(x-1)
y=(4/5)x+31/5